Week of 200203

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Portal
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • local: Julia (WLCG), Kate (WLCG, chair), Sabine (ATLAS), Olga (computing), Borja (monitoring), Aritz (storage)
  • remote: Xavier (KIT), Christoph (CMS), Andrew (TRIUMF), Di (TRIUMF), Elena (CNAF), Darren (RAL), Xin (BNL), Dave (FNAL)

Experiments round table:

  • ATLAS reports ( raw view) -
    • Stable running this week with around 400k job slots reaching 500k with CERN-P1 and NERSC, 25-35 GB/s transfer rate
    • Reprocessing ongoing with data Carousel
    • FTS improved after database improvement and stop of ATLAS data rebalancing for few days
    • Some site disks full waiting rebalancing to restart

  • CMS reports ( raw view) -
    • Many US sites got hit by unlucky IGTF update last week
      • Mainly affected data transfers, but also job submission
      • Fixed after deployment of most recent certificate package
    • Otherwise no major issues

  • ALICE -
    • NTR

  • LHCb reports ( raw view) -
    • Activity:
      • Stripping campaign ongoing
      • Staging will re-start this week based on the stripping activity
    • Issues:
      • no significant issues

Sites / Services round table:

  • ASGC: nc
  • BNL: NTR
  • CNAF: NTR
  • EGI: nc
  • FNAL: dCache tape was moved to 5.2, disk will be moved next week
  • IN2P3: NTR
  • JINR: nc
  • KISTI: NTR
  • KIT: Minor KIT network interruption because of changes to system configuration for integrating new ESX virtualization hosts, this supposedly was not affecting GridKa operations. However, all dCache doors were briefly isolated because of an issue with the high availability setup at about 12 a.m.
  • NDGF: nc
  • NL-T1: SURFsara downtime planned for February 10: https://goc.egi.eu/portal/index.php?Page_Type=Downtime&id=28340
  • NRC-KI: nc
  • OSG: nc
  • PIC: nc
  • RAL: NTR
  • TRIUMF:Large number of files got purged out of tape buffer which cause FTS transfer errors.

  • CERN computing services: NTR
  • CERN storage services:
    • Issue with the performance of FTS during the ATLAS data carousel exercise was fixed by the database on demand team. The FTS team would like to thank the DBoD team for their help and support. A couple of bugs in Rucio were also acknowledged from the Rucio side that prevented to throttle the transfers and actually sent 3 million staging requests to FTS in one go (along with the ongoing re-balancing activity - bringing the number of the requests to around 9 million queued transfers/requests).
  • CERN databases: nc
  • GGUS:
    • A downtime has been scheduled for Wed 06:30-10:30 UTC
      • To move the back-end ARServer from RHEL6 to RHEL7 (it failed last week)
    • Some test alarms may be submitted afterwards
  • Monitoring:
    • Draft availability reports for January send around
  • MW Officer: nc
  • Networks: NTR
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r22 < r21 < r20 < r19 < r18 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r22 - 2020-02-03 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback